OpenStudy (anonymous):

The equation 12x^4 + 20x^3 - 11x^2 - 5x + 2 = 0 has one integer root. Find this root and hence, solve the equation completely. [I found the integer root already that is x + 2, but I do not know how to solve it because I do not know the method in finding out factor that is in fraction. who can teach me with this please..?]

OpenStudy (anonymous):

If "completely" means "find all the roots" you will have to solve a cubic after you have factored out (x+2).

OpenStudy (anonymous):

yes. but I do not know the way in searching out the factors of the cubic, 12x^3 - 4x^2 - 3x + 1. one of its factors is 1/3, the other one is a complex number. do you know the step to find the factor that is a fraction?

OpenStudy (anonymous):

The integer root is -2 and the other roots are 1/2, -1/2, and 1/3.

OpenStudy (anonymous):

nope.. the answers are -2, 1/3 and + or - 1/2

OpenStudy (anonymous):

Use synthetic division, but before you do divide both sides of the equation by 12 so you have the leading coefficient equal to 1.

OpenStudy (anonymous):

hm.. what's synthetic division? After divide both side with 12, then should I find the factor of 1/12?

OpenStudy (anonymous):

i could help too

OpenStudy (anonymous):

now since you know that x=-2 is a root for synthetic division write d coefficients of the equation as 12 20 -11 -5 2

OpenStudy (anonymous):

The synthetic algorithm is fine for the first root because you have been given it but doesn't help for the rest because you don't know (officially) what the roots are, so you have to solve a cubic.

OpenStudy (anonymous):

\[ -2\lfloor 12 20 -11 -5 2 \]

OpenStudy (anonymous):

\[-2\lfloor 12 20 -11 -5 2\]

OpenStudy (anonymous):

-2 | 12 20 -11 -5 2 0

OpenStudy (anonymous):

now write the diff or sum of the first term i.e -2| 12 20 -11 -5 2 | 0 | 12

OpenStudy (anonymous):

multiply d divisor with the remainder and write the no below the second term

OpenStudy (anonymous):

-2| 12 20 -11 -5 2 | 0 24 | 12

OpenStudy (anonymous):

sorry its -24

OpenStudy (anonymous):

as you keep doing this you finally get -2 | 12 20 -11 -5 2 | 0 -24 8 6 -2 | 12 -4 -3 1 0

OpenStudy (anonymous):

As -2 is the root you get d last term as zero, now u hav 4 terms in d bottom most line.. i.e a 3rd order equation

OpenStudy (anonymous):

\[12x ^{3}-4x^{2}-3^{x}+1=0\]

OpenStudy (anonymous):

solve this.. take 4x^2 common in d first 2 terms

OpenStudy (anonymous):

\[4x ^{2}(3x-1)-1(3x-1)=0\]

OpenStudy (anonymous):

now take 3x-1 as common

OpenStudy (anonymous):

\[(3x-1)(4x ^{2}-1)=0\]

OpenStudy (anonymous):

from d first equation you get x=1/3

OpenStudy (anonymous):

from the second you get x=+-1/2